# Self-Classification Enhancement and Correction for Weakly Supervised Object Detection

### v1

@@ 3.3 Self-Classification Enhancement
The SCE module comprises two parallel branches: one for the base multi-class classification and the other for an enhanced intra-class binary classification. These branches work harmoniously during online training, supplementing each other’s strengths.
Multi-Class Classification.
The multi-class classification (MCC) layer shares the same structure with the original OIC layer, containing an FC layer followed by a softmax.
Intra-Class Binary Classification.
We incorporate the intra-class binary classification (ICBC) task to enhance the network’s discrimination between intra-class positive and mis-located samples. To maintain consistency with the MCC layer, we adopt a simple yet effective design to achieve the ICBC task. Specifically, given the proposal feature vector $f$, the ICBC branch consists of an FC layer to generate score matrices and a sigmoid function for normalization:
@@
Sampling strategy for ICBC task.
We adopt MCC results to select training samples $U$ for the ICBC task. Specifically, we first choose a set of top-scoring proposals as the pseudo seed boxes $S=\left\{S_{1},S_{2},\cdots,S_{N}\right\}$ according to MCC results. After that, we propose different strategies to select positive and mis-located samples based on these seed boxes.
@@
However, the samples selected from IoU sampling strategy are insufficient for achieving the ICBC task. Given the limited number of training samples, we further apply a gridding sampling strategy to generate additional mis-located samples for augmentation. Specifically, for each seed box with class $c$, we first apply box scaling by a scaling factor $\theta=0.5$. In this way, the width and height of the seed box are randomly sampled in $[(1-\theta)w,(1+\theta)w]$ and $[(1-\theta)h,(1+\theta)h]$, respectively. The center of the box remains unchanged. Afterward, we generate an $n\times n$ grid on the scaled seed box. We treat each grid as a potential mis-located sample that may only contain a part of an object. It is worth noting that although some grids containing only background may inevitably be selected, their impact is minimal due to their limited quantity. These selected grids $G=\left\{G_{1},G_{2},\cdots,G_{N}\right\}$ are then fed to the RoI pooling layer to generate region features, which are subsequently passed to the ICBC layer. Their pseudo labels are assigned as follows: $y^{\text{icbc}}_{c,G_{i}}=0,G_{i}\in G$.
@@
ICBC Guided Seed Mining.
Seed box mining plays a significant role in training the MCC branch. Some methods pursue the accuracy of seed boxes by using top-scoring strategies Tang et al. (2017), while some others focus on the recall of seed boxes by relaxing the top criteria and adopting non-maximum suppression (NMS) algorithm to remove redundant ones Ren et al. (2020). Different from them, we adopt a soft scoring threshold to mine accurate seed boxes and utilize ICBC results for further fine-tuning.
@@ 3.4 Self-Classification Correction
$IND\leftarrow$$\left\{i|\max_{c}\mathcal{M}_{i,c}>\tau_{midn}\right\}$triangleright confident proposals
$C_{P}\leftarrow$ $\arg\max_{c}$ $\mathcal{M}_{IND,c}$triangleright instance-level classification
$C_{I}\leftarrow$ unique $C_{P}$triangleright image-level classification
$C_{N}\leftarrow\mathcal{C}-C_{I}$triangleright non-exist categories
@@
for $c$ in range ($C_{N}$do
     $\mathcal{F}_{:,c}\leftarrow\lambda\mathcal{F}_{:,c}$triangleright reduce scores
@@
Additionally, to make the SCC algorithm more effective, we adopt the misclassification tolerance (MCT) strategy Wu et al. (2024) to further improve the classification ability of MIDN. The motivation behind the MCT strategy is to introduce tolerance for unrepresentative samples with high semantic similarity to an incorrect class, thereby preventing these samples from dominating the training process and forcing the model to memorize them. Suppose $N_{p}$ classes are present in the image, we identify the misclassified classes containing unrepresentative samples when their score rankings fall within the range of $[N_{p},N_{p}+T_{n}]$, and assign their class weights to $a$ when calculating $L_{MIDN}$. Similarly, the corresponding incorrect classes, whose score rankings fall within the range of $[0,N_{p}]$, are also assigned the same class weight.
@@ 4.2 Implementation Details
[…]
@@ 4.4 Ablation Study
Effect of Each Component.
We present experimental results on VOC 2007 to validate the effectiveness of each component, as summarized in Table 3. Starting with the basic WSOD framework, referred to as the “baseline”, we achieve an initial mAP of 54.6%. Incorporating the ICBC task into the WSOD framework leads to a notable improvement of 0.6%, underscoring the effectiveness of the ICBC branches in enhancing the network’s ability to distinguish between positive and mislocated samples in a class-wise manner. Subsequently, integrating the IGSM algorithm further improves seed box quality, boosting performance to an mAP of 56.7%. Overall, the self-classification enhancement module delivers a significant mAP gain of 2.1%.
@@
[…]
with gridding sampling ($n$=3) $57.9_{\downarrow 0.3}$
@@
Effect of ICBC sampling strategy.
We conduct experiments by employing different sampling strategies for the training of ICBC task, as shown in the upper part of the Table 4. Among all the settings, the combination of the IoU sampling strategy and the gridding strategy achieves the best performance and is insensitive to variations in grid size.
Effect of IGSM algorithm.
We conduct experiments to evaluate each component of IGSM algorithm, as shown in the lower part of the Table 4. On one hand, we first replace our scoring strategy with original top-$1$ strategy, which only selects the proposals with highest scores as seed boxes, resulting in a 1.0% mAP drop, as our scoring strategy more effectively locates various objects using class-wise soft thresholds. Additionally, removing ICBC-guided fine-tuning causes a 1.2% mAP drop, highlighting the superior capability of ICBC in distinguishing positive samples from closely related but mislocated ones.
Effect of Thresholds in IGSM algorithm.
The first two images in Figure 6 illustrate the impact of the scoring threshold $\alpha$ and NMS threshold $\tau_{nms}$ in ICBC guided seed mining (IGSM) algorithm. If the restrictions are wide with low $\alpha$ or high $\tau_{nms}$, more noisy samples will be selected as seed boxes, thus exerting negative impacts on the quality of pseudo labels; if the restrictions are tight with high $\alpha$ or low $\tau_{nms}$, the recall of seed boxes will be reduced, which hinders the improvement brought from the IGSM algorithm. Compared with the baseline, our IGSM is not sensitive to the choice of values around the optimal ones ($\alpha=0.9,\tau_{nms}=0.1$), and consistently delivers at least 2.8% mAP.
Effect of Scoring Factor in SCC algorithm.
The last two images in Figure 6 illustrate the impact of the scoring factor $\lambda$ and the scoring threshold $\tau_{midn}$ in the Self-Classification Correction (SCC) algorithm. SCC algorithm effectively reduces the mis-classified samples by predicting the non-exist categories and reducing their scores, thus improving the detection performance. In general, the performance is not sensitive to the choice of values around the optimal ones ($\lambda=0.01,\tau_{midn}=0.001$) when the $\tau_{midn}$ is effective ($\tau_{midn}>0$), with the highest gap no more than 0.2% mAP. We attribute it to that, SCC algorithm effectively reduces the mis-classified samples by predicting the non-exist categories and reducing their scores, thus improving the detection performance.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants (62422204, 62472139) and in part by the Zhejiang Provincial Natural Science Foundation of China under Grant LQN25F030014, LDT23F02025F02. This work is also supported by the Open Project Program of the State Key Laboratory of CAD&CG (Grant No. A2403), Zhejiang University.

### v2

@@ 3.3 Self-Classification Enhancement
The SCE module comprises two parallel branches: one for the base multi-class classification and the other for an enhanced intra-class binary classification. These branches work harmoniously during online training, supplementing each other’s strengths. Multi-Class Classification. The multi-class classification (MCC) layer shares the same structure with the original OIC layer, containing an FC layer followed by a softmax.
Intra-Class Binary Classification. We incorporate the intra-class binary classification (ICBC) task to enhance the network’s discrimination between intra-class positive and mis-located samples. To maintain consistency with the MCC layer, we adopt a simple yet effective design to achieve the ICBC task. Specifically, given the proposal feature vector $f$, the ICBC branch consists of an FC layer to generate score matrices and a sigmoid function for normalization:
@@
Sampling strategy for ICBC task. We adopt MCC results to select training samples $U$ for the ICBC task. Specifically, we first choose a set of top-scoring proposals as the pseudo seed boxes $S=\left\{S_{1},S_{2},\cdots,S_{N}\right\}$ according to MCC results. After that, we propose different strategies to select positive and mis-located samples based on these seed boxes.
@@
However, the samples selected from IoU sampling strategy are insufficient for achieving the ICBC task. Given the limited number of training samples, we further apply a gridding sampling strategy to generate additional mis-located samples for augmentation. Specifically, for each seed box with class $c$, we first apply box scaling by a scaling factor $\theta=0.5$. In this way, the width and height of the seed box are randomly sampled in $(1-\theta)w,(1+\theta)w]$ and $(1-\theta)h,(1+\theta)h]$, respectively. The center of the box remains unchanged. Afterward, we generate an $n\times n$ grid on the scaled seed box. We treat each grid as a potential mis-located sample that may only contain a part of an object. It is worth noting that although some grids containing only background may inevitably be selected, their impact is minimal due to their limited quantity. These selected grids $G=\left\{G_{1},G_{2},\cdots,G_{N}\right\}$ are then fed to the RoI pooling layer to generate region features, which are subsequently passed to the ICBC layer. Their pseudo labels are assigned as follows: $y^{\text{icbc}}_{c,G_{i}}=0,G_{i}\in G$.
@@
ICBC Guided Seed Mining. Seed box mining plays a significant role in training the MCC branch. Some methods pursue the accuracy of seed boxes by using top-scoring strategies Tang et al. (2017), while some others focus on the recall of seed boxes by relaxing the top criteria and adopting non-maximum suppression (NMS) algorithm to remove redundant ones Ren et al. (2020). Different from them, we adopt a soft scoring threshold to mine accurate seed boxes and utilize ICBC results for further fine-tuning.
@@ 3.4 Self-Classification Correction
$IND\leftarrow$ $\left\{i|\max_{c}\mathcal{M}_{i,c}>\tau_{midn}\right\}$ $\triangleright$ confident proposals
$C_{P}\leftarrow$ $\arg\max_{c}$ $\mathcal{M}_{IND,c}$ $\triangleright$ instance-level classification
$C_{I}\leftarrow$ unique $C_{P}$ $\triangleright$ image-level classification
$C_{N}\leftarrow\mathcal{C}-C_{I}$ $\triangleright$ non-exist categories
@@
for $c$ in range ($C_{N}$) do
  $\mathcal{F}_{:,c}\leftarrow\lambda\mathcal{F}_{:,c}$ $\triangleright$ reduce scores
@@
Additionally, to make the SCC algorithm more effective, we adopt the misclassification tolerance (MCT) strategy Wu et al. (2024) to further improve the classification ability of MIDN. The motivation behind the MCT strategy is to introduce tolerance for unrepresentative samples with high semantic similarity to an incorrect class, thereby preventing these samples from dominating the training process and forcing the model to memorize them. Suppose $N_{p}$ classes are present in the image, we identify the misclassified classes containing unrepresentative samples when their score rankings fall within the range of $N_{p},N_{p}+T_{n}]$, and assign their class weights to $a$ when calculating $L_{MIDN}$. Similarly, the corresponding incorrect classes, whose score rankings fall within the range of $0,N_{p}]$, are also assigned the same class weight.
@@ 4.2 Implementation Details
Following a widely-used setting, we adopt VGG16 Simonyan and Zisserman (2014) pre-trained on ImageNet Deng et al. (2009) as the backbone and Selective Search Uijlings et al. (2013) for proposal generation. The whole framework is end-to-end optimized using stochastic gradient descent (SGD), and the momentum, weight decay, and batch size are set as 0.9, $5\times 10^{-4}$, and 4, respectively. The initial learning rate is set as $1\times 10^{-3}$ for the first 70kk, 170kk iterations, and it is dropped by a factor of 10 for the following 20kk, 40kk iterations for VOC 2007 and VOC 2012, respectively. We set $\alpha$ to 0.9 and the NMS threshold $\tau_{nms}$ to 0.1 in the SEC module. $\lambda$ and $\tau_{midn}$ in SCC algorithm are set to 0.01 and 0.001, respectively. The hyperparameters of MCT strategy are set as the same with Wu et al. (2024), i.e., $T_{n}=1,a=0.4$. The loss weight $\gamma$ is set to 0.1 for the training balance. Following the previous WSOD works, $\tau_{l}$, $\tau_{h}$, and $T$ are set to 0.1, 0.5, and 3, and the images are multi-scaled to {480, 576, 688, 864, 1000, 1200} for both training and inference.
@@ 4.4 Ablation Study
Effect of Each Component. We present experimental results on VOC 2007 to validate the effectiveness of each component, as summarized in Table 3. Starting with the basic WSOD framework, referred to as the “baseline”, we achieve an initial mAP of 54.6%. Incorporating the ICBC task into the WSOD framework leads to a notable improvement of 0.6%, underscoring the effectiveness of the ICBC branches in enhancing the network’s ability to distinguish between positive and mislocated samples in a class-wise manner. Subsequently, integrating the IGSM algorithm further improves seed box quality, boosting performance to an mAP of 56.7%. Overall, the self-classification enhancement module delivers a significant mAP gain of 2.1%.
@@
with gridding sampling (nn=2) $58.2_{\ 0.0}$
with gridding sampling (nn=3) $57.9_{\downarrow 0.3}$
@@
Effect of ICBC sampling strategy. We conduct experiments by employing different sampling strategies for the training of ICBC task, as shown in the upper part of the Table 4. Among all the settings, the combination of the IoU sampling strategy and the gridding strategy achieves the best performance and is insensitive to variations in grid size.
Effect of IGSM algorithm. We conduct experiments to evaluate each component of IGSM algorithm, as shown in the lower part of the Table 4. On one hand, we first replace our scoring strategy with original top-$1$ strategy, which only selects the proposals with highest scores as seed boxes, resulting in a 1.0% mAP drop, as our scoring strategy more effectively locates various objects using class-wise soft thresholds. Additionally, removing ICBC-guided fine-tuning causes a 1.2% mAP drop, highlighting the superior capability of ICBC in distinguishing positive samples from closely related but mislocated ones.
Effect of Thresholds in IGSM algorithm. The first two images in Figure 6 illustrate the impact of the scoring threshold $\alpha$ and NMS threshold $\tau_{nms}$ in ICBC guided seed mining (IGSM) algorithm. If the restrictions are wide with low $\alpha$ or high $\tau_{nms}$, more noisy samples will be selected as seed boxes, thus exerting negative impacts on the quality of pseudo labels; if the restrictions are tight with high $\alpha$ or low $\tau_{nms}$, the recall of seed boxes will be reduced, which hinders the improvement brought from the IGSM algorithm. Compared with the baseline, our IGSM is not sensitive to the choice of values around the optimal ones ($\alpha=0.9,\tau_{nms}=0.1$), and consistently delivers at least 2.8% mAP.
Effect of Scoring Factor in SCC algorithm. The last two images in Figure 6 illustrate the impact of the scoring factor $\lambda$ and the scoring threshold $\tau_{midn}$ in the Self-Classification Correction (SCC) algorithm. SCC algorithm effectively reduces the mis-classified samples by predicting the non-exist categories and reducing their scores, thus improving the detection performance. In general, the performance is not sensitive to the choice of values around the optimal ones ($\lambda=0.01,\tau_{midn}=0.001$) when the $\tau_{midn}$ is effective ($\tau_{midn}>0$), with the highest gap no more than 0.2% mAP. We attribute it to that, SCC algorithm effectively reduces the mis-classified samples by predicting the non-exist categories and reducing their scores, thus improving the detection performance.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants (62422204, 62472139), in part by the Fundamental Research Funds for the Provincial Universities of Zhejiang under Grant No. GK259909299001-040, and in part by the Zhejiang Provincial Natural Science Foundation of China under Grant LQN25F030014, LDT23F02025F02. This work is also supported by the Open Project Program of the State Key Laboratory of CAD&CG (Grant No. A2403), Zhejiang University.